[figure: 9.7 cm]

MAGNETORESISTANCE IN SEMICONDUCTOR-FERRITE

STRUCTURE

A.S. Bajtsar^{∗}, R.I. Baitsar, S.S. Varshava, S.I. Yushchuk ∗Lviv I.Franko State University, 8 Kyryla and Mefodia str., Lviv, 290005, Ukraine

State University ”Lviv Politechnica”, Bandery str.12, Lviv, 290646, Ukraine

Semiconductor-ferrite structures, formed by means of tight contact between InSb monocrystal (magnetoresister–MR) and ferrite substrate have been investigated in this work. The oriented Mn_{x}Zn_{1−x}F e_{2}O_{4} monocrystal substrates, Mn-Zn polycrystal ferrite and YIG films epitaxially grown on GGG surface from liquid-phase (LPE) were used. The substrate influence on the VAC and the ΔU (ΔR) increases depending on the magnetic field (B) for different I_{0}. Substrate magnetization direction is defined by the anisotropy of crystal growth. Approximatly two fold increase of susceptibility depending on the field γ=157%/Tl (0-0.15 Tl) and 64%/Tl (0.15-0.5 Tl) at I_{0}=160 mA was obtained. The outward magnetic field screening was achieved by placing MR in the ferrite Mn-Zn grooves. As a result the ΔU independence of the outward field and N-like ΔU=f(B) characteristic were observed.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : R.I. Baitsar

Address for correspondence : State University ”Lviv Politechnica”, Bandery str.12, Lviv, 290646, Ukraine

Email address : iostrov@polynet.lviv.ua